OpenStudy (anonymous):

simplify by removing factors of 1:y^2-11y+18/y^2-7y-18

OpenStudy (anonymous):

y^2-11y+18 = (y-9) (y-2) y^2-7y-18 = (y - 9) ( y + 2) answer is \[\frac{y-2}{y+2}\]